HSO Report – December 26th

category operations-report]

Crew 236 Operations Report 26-DEC-2021

SOL: 7

Name of person filing report: Pavithra Ravi

Non-nominal systems: None

ROVERS

Spirit rover used: No

Hours (before EVA): N/A

Beginning charge (Before EVA): N/A

Ending charge (On return from EVA, before recharging): N/A

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: Yes

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

General notes and comments: The front airlock door has been secured
with the rope that was attached to it, to keep it closed during the
high winds. The back airlock has been kept closed by lodging a wrench
in the locking mechanism. The outer surface of the panels near the
back airlock area (next to the static tank) came loose during the
winds. Thankfully, there is no breach. There are some materials strewn
through the tunnels as a result of the high winds. We can provide
pictures if needed. I believe Shannon has already inspected it.

Summary of Hab operations: Nominal

WATER USE:

Water (static tank): approx. 265 gallons

Water (loft tank): 54 gallons

Water Meter: 01540912 units

Static to Loft Pump used – yes or no: Yes. It was filled from 17 to 54
gallons at about 20:00 today.

Static tank pipe heater (on or off): On

Static tank heater (on or off) Off

Toilet tank emptied (no or yes): Yes, at 9:30 PM last night

Summary of internet: Nominal

Summary of suits and radios: Nominal.

Summary of GreenHab operations: Nominal

WATER USE: 12.5 gallons were used today. This leaves about 282
gallons in the tank.

Heater (on or off): On

Supplemental light (on or off): On

Harvest: None

Summary of ScienceDome operations: Nominal

Dual split, Heat or AC (on or off): off

Summary of RAM operations: Tools used, work done

Some zipties and a wrench were obtained to keep the back airlock door
closed during the high winds. The wrench was ultimately jammed in the
lock to keep the door securely closed. The zipties were unused and
will be returned.

Summary of any observatory issues: Nothing new to report.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support:

None

GreenHab Officer – December 24TH

[category: GreenHab-Report]

Crew 236 GreenHab Report 12242021
GreenHab Officer: Vladimir Zeltsman
Environmental Control: Around 17% Humidity
Average Temperatures: 70 degrees
Hours of Supplemental Light: 7 hours
Daily Water Usage For Crops: 9.5 gallons
Daily Water Usage For Research and/or Other Purposes: None
Water In Blue Tank: 203.5 Gallons
Times of watering for crops: 11:30 am
Changes to Crops: None
Narrative: None
Harvest: 2g Basil, 37g Tomato, 1g Green Onion
Support/supplies needed: None

GreenHab Officer – December 24TH

MJ!

Sorry to reply so late, we’ve been fixing a watery situation at the outpost.
I think you might have the wrong email to send the reports. Gimme a second

On Fri, Dec 24, 2021, 19:23 MJ Marggraff <mj.red.planet.rising> wrote:

[category: GreenHab-Report]

Crew 236 GreenHab Report 12242021
GreenHab Officer: Vladimir Zeltsman
Environmental Control: Around 17% Humidity
Average Temperatures: 70 degrees
Hours of Supplemental Light: 7 hours
Daily Water Usage For Crops: 9.5 gallons
Daily Water Usage For Research and/or Other Purposes: None
Water In Blue Tank: 203.5 Gallons
Times of watering for crops: 11:30 am
Changes to Crops: None
Narrative: None
Harvest: 2g Basil, 37g Tomato, 1g Green Onion
Support/supplies needed: None

GreenHab Officer – December 24TH

[category: GreenHab-Report]

Crew 236 GreenHab Report 12242021
GreenHab Officer: Vladimir Zeltsman
Environmental Control: Around 17% Humidity
Average Temperatures: 70 degrees
Hours of Supplemental Light: 7 hours
Daily Water Usage For Crops: 9.5 gallons
Daily Water Usage For Research and/or Other Purposes: None
Water In Blue Tank: 203.5 Gallons
Times of watering for crops: 11:30 am
Changes to Crops: None
Narrative: None
Harvest: 2g Basil, 37g Tomato, 1g Green Onion
Support/supplies needed: None

Crew 260 Sol Summary for November 2nd

[category Sol Summary – report}

Crew 260 Sol Summary Report 02-11-2021

Sol: 1

Summary Title: I got a rock!

Author’s name: Jen Carver-Hunter

Mission Status: Nominal

Sol Activity Summary: Our morning started with a lesson about Mars’ geologic periods and analogous processes that occurred at MDRS, forming the specimens we decided to collect on EVAs. Both EVA teams successfully completed their missions. After successful EVAs, crew members identified and labeled their samples, which they added to their “Mars rocks” kits.

Look Ahead Plan: Finish sample collection (EVA 3), sample categorization and analysis, rocket construction

Anomalies in work: No anomalies

Weather: Sunny, warm, and pleasant

Crew Physical Status: One crew member fell down a few stairs during the night and injured their knee. Injury was not serious enough to need professional medical attention, but has limited crewmember’s ability to participate in EVAs. The rest of the crew has made modifications to support all crew members while maintaining sim.

EVA: Geologic sample collection on the road to Barranca Butte and Kissing Camel Ridge.

Reports to be filed: Operations report, Journalist report, EVA request

Support Requested: None at this time.

Operations Report – September 16th

[category operations-report]

MDRS Complete Operations Report 16-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine daily

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 92.8

Minimum 80

Maximum 100

Note on solar: Nothing to report

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 80%

Propane Reading, director tank— 74%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (although I can’t find the gas can)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 225 gallons

Static to Loft Pump used – no

Water in GreenHab – 20 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 210 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours: 219.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours: 136.9

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: 72.3

Beginning charge: 100

Ending charge: 95

Currently charging: no

Curiosity rover used: no

Hours: 144.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: I was looking at the problem with the rovers regarding the damage they do to the spacesuits. What really needs to happen is to attach the seat back in the manner that we did to the former Deimos, now Percy.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no

Reason for use: nothing to report

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to gather garbage to take to town tomorrow. I forgot to note earlier that the shifter has broken but it is still usable.

CrewCar used and why, where? Yes, to town to do laundry. I forgot to note yesterday that I drove it on the dry, paved roads in 4×4. It had been put in 4×4 by someone on the work party, and I didn’t notice it before I drove it. I tested it in all 4 modes and there appears to be no damage. I do need to get this vehicle in for repairs as soon as I can get someone to drive with me to drop it off for service. It needs work on the undercarriage.

Luna used and why, where? Yes, to go to storage units and pick up all materials related to spacesuits

General notes and comments: The storage units are not yet cleaned out but they have been emptied of much unused materials. What will be left when completed will be stored in a single unit and will include the HabCar rear seat, rear seat backs for the rovers, Ikea materials for the ScienceDome and the observatory materials. The power tools will be relocated to the station, and what little remains will either be stored here in the shed, given away, or stored with the URC materials in the one remaining unit.

Summary of internet: All nominal.

Summary of suits and radios: The NorCal Chapter representatives will be here tomorrow with the suits. In addition, while they were here they will go through the suit materials that have been stored in the storage unit. I brought these materials out to the station today so they will be able to go through them here.

Campus wide inspection, if action taken, what and why? Lift boom (or is it boom lift?) and power washer are still here. Not sure when they will be picked up.

Summary of general operations: Nothing to report

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Replaced the clothesline (such excitement!)

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing to report

Operations Report – April 17th

[category operations-report]

Crew 245 Operations Report 17-04-2021

SOL: 7

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7

From what time last night: 10:00 pm, SOC 57%

To what time this morning: 5:00 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 45% at 7:40 pm (Before generator is run at night),

Diesel Reading – 50%

Station Propane Reading – 73%

Water (loft tank): 15 gallons

Water Meter: 150460.4 units

Water (static tank): 340 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: Yes

Hours: 203.3

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: 350.2

Reason for use: EVA

Oil Added? No

ATV Fuel Used: 1 gallon (estimated)

# Hours the ATVs were used today: 6 (but 1.5 hours of active transit)

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission

HabCar used and why, where?: No

CrewCar used and why, where?: No

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew did not consume allotted bandwidth last night so there was an additional bandwidth bonus available today.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 3, 4, 6, 7, 9 and 11 all used for EVA today. Suit 7 battery would not charge after EVA, and Suit 9 battery was not producing sufficient power output even when fully charged, so both were replaced.

Summary of Hab operations: Nothing to report besides nominal crew activities.

Summary of GreenHab operations: Being used to host Crew Botanist botany experiment. Crew Botanist briefly checked plant growth and health in GreenHab today.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report.

Final Schedule for the 2019-2020 Field Season

This one can be released to the public:

Final Schedule for the 2019-2020  field season:

Oct 17-20, 2019—Mars Society Conference, Los Angeles

Oct 22-25— IKEA Work Party

Crew 214—Oct 26- Nov10— MSA ExBoomerang

Crew 215—Nov 9-24, 2019— MSA ExBoomerang

Crew 216—Nov 23—Dec 8, 2019— A.R.E.S. (Mars Society crew)

Crew 217—Dec 7-14— Spaceward Bound Utah (Utah NASA Space Consortium Grant)

Dec 14-22— Holiday Break

Crew 218—Dec 21, 2019- Jan 5, 2020— Purdue

Crew 219—Jan 4-19 — ARES  (Florida Tech)

Crew 220—Jan 18-Feb 2— Mars Academy

Crew 221—Feb 1-16— Earth 2 Mars

Crew 222—Feb 15-Mar 1 —Supaero I

Crew 223—Feb 29-Mar 15— Supaero II

Crew 224—Mar 14-21—Wilderness Medical Society class

Crew 225—Mar 22-29—University of Colorado 2 Mars class

Crew 226—Mar 28-Apr 12–  Colombia

Crew 227—Apr 11-26—UCL

Crew 228—Apr 25-May 10—Mars Society II

Crew 229—May 9-24—Work Party

May 28-30– University Rover Challenge

Crew 230– June 1-8, 2020— Medical Makers

Crew 231—June 8-15—Martian Biology II

June 22-28—Astronomy Team Summer Residency